Evaluate the Mersenne number M17, and determine whether it is prime.?
Answer #1

M17 = 2 - 1 = 131072-1 = 131071 Mersenne number 17 is of the form Mn = 2h - 1 so, L-e M17 = 131071 Since factor of 131071 is